OpenStudy (anonymous):

Simplify (2√5+3√7)^2. show your work, justify each step

OpenStudy (pfenn1):

\[\left( 2\sqrt{5}+3\sqrt{7} \right)^{2}=\] multiply out\[=(2\sqrt{5})^{2}+2(2\sqrt{5})(3\sqrt{7})+(3\sqrt{7})^{2}=\]

OpenStudy (anonymous):

thank you u r awesome!!!

OpenStudy (pfenn1):

you're welcome.

OpenStudy (pfenn1):

You do know you can simplify further? For example, \[(2\sqrt{5})^{2}=4 (\sqrt{5})^{2}=4 \times 5 = 20\]

OpenStudy (anonymous):

I think i have it now

OpenStudy (pfenn1):

great!